Jill's present age is three times her sister's age two years ago. If Jill is 18 years old, how old is her sister now?
Found 2 solutions by checkley77, Ruli:
Answer by checkley77(12844)

18=3(X-2)
18=3X-6
3X=18+6
3X=24
X=24/3
X=8 YEARS OLD IS THE SISTER NOW.

Answer by Ruli(21)

Her sister's age 2 years ago = (x - 2)
Jill's present age = 3(x - 2)=18
3(x - 2) = 18
3x - 6 = 18
3x - 6 + 6 = 18 + 6
3x = 24
x = 8
check:
3(8 - 2)
3(6)= 18